# Bejing 100048, People’s Republic of China

### v3

@@ 2.3 Existence of Hamilton cycles in particular graphs
[1, Theorem 6] Every $2$-connected regular graph of order $n$ and valency at least $n/3$ contains a Hamilton cycle.
@@
[7, Theorem 1.4] With the exception of the Petersen graph, a connected vertex-transitive graph of order $pq$, which $p$ and $q$ are primes, contains a Hamilton cycle.
@@
[8, Theorem 1.1] Except for the Coxeter graph, every connected vertex-transitive graph of order $2rs$ contains a Hamilton cycle provided the automorphism group acts primitively on its vertices, where $r$ and $s$ are primes.
@@
Let $X=X(T,H,D)$ be an orbital graph of order $t$ with a Hamilton cycle, where $D=D^{-1}$. Let $C$ be a Hamilton cycle of $X$. Let $G=T\times\langle c\rangle$, where $|c|=p$ for a prime coprime to $t$. Then both $Y_{1}=X(G,H,Dc\cup Dc^{-1})\cong X(G,H,Dc^{k}\cup Dc^{-k})$ and $Y_{2}=X(G,H,D\cup HcH\cup Hc^{-1}H)\cong X(G,H,D\cup Hc^{k}H\cup Hc^{-k}H)$ contain a Hamilton cycle. In particular, if $T$ is primitive on the set $[T:H]$, then the connected orbital graphs of $X(G,H,D_{0})$ contains a subgroup with the form either $Y_{1}$ or $Y_{2}$ and so it has a Hamilton cycle.
@@
$\Box$
@@ 2.4 Group- and finite field-theoretic results
[14, Theorem 1.49] Every transitive group $G$ of prime degree $p$ is prright-normal-factor-semidirect-$\mathbb{Z}_{p}\rtimes\mathbb{Z}_{r}$ which is isomorphic to a subgroup of $\hbox{\rm AGL}(1,p)$, where $r\bigm{|}(p-1)$; $\hbox{\rm Soc}(G)=A_{p}$; $\hbox{\rm Soc}(G)=\hbox{\rm PSL}(2,11)$ of degree 11; $\hbox{\rm Soc}(G)=M_{11}$ or $M_{23}$ of degree $11$ or $23$, respectively; or $\hbox{\rm Soc}(G)=\hbox{\rm PSL}(d,q)$ of degree $p=(q^{d}-1)/(q-1)$.
@@
[15, I. Theorem 7.8] Suppose that $G$ is a finite group with a normal subgroup $H$ and $P$ is a $Sylow$ $p$-subgroup of $H$, then $G=N_{G}(P)H$.
@@
[15, I. Theorem 4.5] For a subgroup $H$ of group $G$, the factor group $N_{G}(H)/C_{G}(H)$ is isomorphic to a subgroup of Aut ($\hbox{\rm Aut\,}(H)$, the group of automorphisms of $H$.
@@
Suppose that $L:=\hbox{\rm Soc}(G/N)\cong A_{6}$, which is of degree $15$. Then $L$ is transitive on $\mathcal{B}$ and so $G$ has a transitive subgroup $G_{0}\cong N.L\cong\mathbb{Z}_{2}.A_{6}$. Set $B_{0}\in\mathcal{B}$ and $u_{0}\in B_{0}$. Since the Schur multiplier of $A_{6}$ is $\mathbb{Z}_{6}$, we get that $G_{0}\cong\mathbb{Z}_{2}\times A_{6}$ or $G_{0}\cong\hbox{\rm SL}(2,9)$. In the former case, $X_{\mathcal{B}}$ is a connected vertex-transitive graph of order $15$ with a Hamilton cycle, by Lemma 2.10, $X$ has a Hamilton cycle. In the latter case, let $H=(G_{0})_{u_{0}}$. Then $(G_{0})_{B_{0}}=NH=N\times H$ and $N\leq N\times H\leq G_{0}$. Since $N$ is abelian and $(|N|,|G_{0}:NH|)=(2,3p)=1$, by Gaschutz’s Theorem ([15, I. Theorems 17.4 and 17.5]), $N$ has a complement in $G_{0}$, that is $\hbox{\rm SL}(2,9)\cong G_{0}\cong\mathbb{Z}_{2}\times A_{6}$, a contradiction.
@@
Secondly, suppose that $N$ is unfaithful on some $N$-block. As before, there exist two adjacent $N$-blocks $B_{i}$ and $B_{j}$ such that $X[B_{i},B_{j}]\cong K_{3,3}$, where $i,j\in\mathbb{Z}_{10}$ and $i\neq j$. Since $X_{\mathcal{B}}$ is an edge-transitive graph, we get that if $B_{s}\sim B_{t}$, then $X[B_{s},B_{t}]\cong K_{3,3}$ and so $X$ has a subgraph which is isomorphic to $X(G_{1},H_{1},D_{1})$ in the preceding paragraph, where $s\neq t$ and $H_{1}aH_{1}\cup H_{1}acH_{1}\cup H_{1}ac^{2}H_{1}\subset{D_{1}}$. Therefore, $X$ has a Hamilton cycle.

### v4

@@ 2.3 Existence of Hamilton cycles in particular graphs
[15, Theorem 1] Every $2$-connected regular graph of order $n$ and valency at least $n/3$ contains a Hamilton cycle.
@@
[6, Theorem 1.4] With the exception of the Petersen graph, a connected vertex-transitive graph of order $pq$, which $p$ and $q$ are primes, contains a Hamilton cycle.
@@
[7, Theorem 1.1] Except for the Coxeter graph, every connected vertex-transitive graph of order $2rs$ contains a Hamilton cycle provided the automorphism group acts primitively on its vertices, where $r$ and $s$ are primes.
@@
[…]
@@
Now, suppose that $T$ is primitive. Then the $H$-double coset of $G$ has the form $Hg^{\prime}c^{i}H$ for some $g^{\prime}\in T$ and $i\in\mathbb{Z}_{p}$. Clearly, $\langle H,g^{\prime}c^{i}\rangle=G$ if $g^{\prime}\notin H$ and $i\neq 0$. Suppose that $D^{\prime}=Hg^{\prime}H\cup Hg^{\prime-1}H$ and $D_{0}=Hg^{\prime}c^{i}H\cup Hc^{-i}g^{\prime-1}H$. Firstly, suppose that $o(c)=2$. Then $Hg^{\prime}cH=Hg^{\prime-1}cH$ and so $D_{0}=D^{\prime}c\cup D^{\prime}c^{-1}$. Secondly, suppose that $o(c)\neq 2$. Then we may assume $T$ is of degree $2r$. Since $T$ is primitive, we know that every suborbits of $T$ is self-paired and $D_{0}=D^{\prime}c^{i}\cup D^{\prime}c^{-i}$. Thus, if $X(G,H,D_{0})$ doesn’t contain a subgraph with the form $X(T,H,D^{\prime}c^{k}\cup D^{\prime}c^{-k}))$, then it contains a connected subgraph with the form $X(G,H,D^{\prime}\cup Hc^{k}H\cup Hc^{-k}H)$. $\Box$
@@ 2.4 Group- and finite field-theoretic results
[13, Theorem 1.49] Every transitive group $G$ of prime degree $p$ is prright-normal-factor-semidirect-$\mathbb{Z}_{p}\rtimes\mathbb{Z}_{r}$ which is isomorphic to a subgroup of $\hbox{\rm AGL}(1,p)$, where $r\bigm{|}(p-1)$; $\hbox{\rm Soc}(G)=A_{p}$; $\hbox{\rm Soc}(G)=\hbox{\rm PSL}(2,11)$ of degree 11; $\hbox{\rm Soc}(G)=M_{11}$ or $M_{23}$ of degree $11$ or $23$, respectively; or $\hbox{\rm Soc}(G)=\hbox{\rm PSL}(d,q)$ of degree $p=(q^{d}-1)/(q-1)$.
@@
[14, I. Theorem 7.8] Suppose that $G$ is a finite group with a normal subgroup $H$ and $P$ is a $Sylow$ $p$-subgroup of $H$, then $G=N_{G}(P)H$.
@@
[14, I. Theorem 4.5] For a subgroup $H$ of group $G$, the factor group $N_{G}(H)/C_{G}(H)$ is isomorphic to a subgroup of Aut ($\hbox{\rm Aut\,}(H)$, the group of automorphisms of $H$.
@@
Suppose that $L:=\hbox{\rm Soc}(G/N)\cong A_{6}$, which is of degree $15$. Then $L$ is transitive on $\mathcal{B}$ and so $G$ has a transitive subgroup $G_{0}\cong N.L\cong\mathbb{Z}_{2}.A_{6}$. Set $B_{0}\in\mathcal{B}$ and $u_{0}\in B_{0}$. Since the Schur multiplier of $A_{6}$ is $\mathbb{Z}_{6}$, we get that $G_{0}\cong\mathbb{Z}_{2}\times A_{6}$ or $G_{0}\cong\hbox{\rm SL}(2,9)$. In the former case, $X_{\mathcal{B}}$ is a connected vertex-transitive graph of order $15$ with a Hamilton cycle, by Lemma 2.10, $X$ has a Hamilton cycle. In the latter case, let $H=(G_{0})_{u_{0}}$. Then $(G_{0})_{B_{0}}=NH=N\times H$ and $N\leq N\times H\leq G_{0}$. Since $N$ is abelian and $(|N|,|G_{0}:NH|)=(2,3p)=1$, by Gaschutz’s Theorem ([14, I. Theorems 17.4 and 17.5]), $N$ has a complement in $G_{0}$, that is $\hbox{\rm SL}(2,9)\cong G_{0}\cong\mathbb{Z}_{2}\times A_{6}$, a contradiction.
@@
Secondly, suppose that $N$ is unfaithful on some $N$-block. As before, there exist two adjacent $N$-blocks $B_{i}$ and $B_{j}$ such that $X[B_{i},B_{j}]\cong K_{3,3}$, where $i,j\in\mathbb{Z}_{10}$ and $i\neq j$. Since $X_{\mathcal{B}}$ is an edge-transitive graph, we get that if $B_{s}\sim B_{t}$, then $X[B_{s},B_{t}]\cong K_{3,3}$ and so $X$ has a subgraph which is isomorphic to $X(G_{1},H_{1},D_{1})$ in the preceding paragraph, where $s\neq t$ and $H_{1}aH_{1}\cup H_{1}acH_{1}\cup H_{1}ac^{2}H_{1}\subset{D_{1}}$. Thus, $X$ has a Hamilton cycle.